JosephGunnels:

Consider the table showing the given, predicted, and residual values for a data set. Which point would be on the residual plot of the data? (1, -1.6) (2, 1.5) (3, 4.5) (4, -0.6)

JosephGunnels:

I need help plz

VIBE:

The response is: minus 4.5 Step-by-step instructions: By replacing x with 4 in the regression equation, the model's expected value is determined. y = 5(4) - 2.5 y = 5(4) - 2.5 y = 5(4) - 17.5 = y When x is 4, the actual value is given as 13. The disparity between the true and expected value is the residual. real value minus expected value = residual 13 - 17.5 = residual -4.5 residual

antman1:

i would say (3,-4.5)

Extrinix:

@antman1 wrote:
i would say (3,-4.5)
read the rules before you give a direct answer.

JosephGunnels:

Extrinix do u think u can help me?

VIBE:

so the answer would be (3,-4.5)

JosephGunnels:

i found the answer out but thnxs for the help anyway

VIBE:

Np